# MaxSup: Overcoming Representation Collapse in Label Smoothing

### v2

@@ Appendix H Extended Comparison to More Label Smoothing Alternatives
Baseline 76.41±0.10
Label Smoothing 76.91±0.11
[…]
$\text{OLS}^{*}$ 77.23±0.21
MaxSup 77.69±0.07
Logit Penalty 76.73±0.10
Confidence Penalty 76.58±0.12
Adaptive Label Smoothing 77.16±0.15
Appendix I Classification on Fine-Grained Datasets
We have further compared MaxSup and other methods on CUB-200 and Stanford Cars for fine-grained classification in Table 11. The performance of MaxSup again supports its superiority against existing approaches.
Table 11: Classification on CUB and Cars Datasets.
Method CUB[34] Cars[14]
no_ls 80.88 90.27
ls 81.96 91.64
ols 82.33 91.96
zipfs 81.40 90.99
max_sup 82.53 92.25
@@ Appendix I Classification on Fine-Grained Datasets
To illustrate the differences between Max Suppression Regularization and Label Smoothing, we follow the projection technique of Müller et al. [21]. Specifically, we select three semantically related classes and construct an orthonormal basis for the plane intersecting their class templates in feature space. We then project each sample’s penultimate-layer activation vector onto this plane. To ensure the visual clarity of the resulting plots, we randomly sample 80 images from the training or validation set for each of the three classes.
@@
    Confusion. We identify a class that the Label Smoothing (LS)–trained model frequently misclassifies and select two additional classes involved in those misclassifications (Figure 3(c), Figure 4(c)). Conversely, we also examine a scenario where a class under Max Suppression is confused with others, highlighting key differences (Figure 3(d), Figure 4(d)).
@@
As shown in Figures 4 and 3, models trained with Max Suppression exhibit:
@@
Appendix K Evaluation of Calibration Performance
We have additionally evaluated the effect of MaxSup on calibration in Table 13, based on the Expected Calibration Error (ECE) [8].
Table 13: Expected Calibration Error (ECE) on CIFAR100. Lower is better.
ce $10.3\%$
ls $3.8\%$
ols $2.9\%$
zipf $3.5\%$
maxsup $2.6\%$
Appendix L Ablation on the Weight Schedule
We conducted an ablation study on the $\alpha$ schedule, as shown in Table 14. The consistently high accuracy across settings demonstrates the robustness of MaxSup. The adaptive $\alpha$ schedule, adopted from [18], further highlights the method’s integrity and compatibility with principled design choices.
@@
Appendix M Analysis of Computation Efficiency
@@
Since $K$ (the number of classes) is usually small (e.g., 1000 for ImageNet-1K), this overhead is minimal compared to the overall forward/backward pass of a deep network. In Table 15, we report the average training time per epoch using a ResNet-50 model on the ImageNet-1K dataset.
@@
Appendix N Logits Visualization
As mentioned in Section 4.1.2, all Label Smoothing variants apply a different penalty on the logits. To illustrate the impact of different methods on the logits, we plot the histogram of logits of ResNet-50 networks trained with each method over the ImageNet validation set, as shown in Figure 5.

### v3

@@ Appendix I Extended Comparison to More Label Smoothing Alternatives
Baseline $76.41\pm 0.10$
Label Smoothing $76.91\pm 0.11$
Zipf-LS $76.73\pm 0.17$
OLS $77.23\pm 0.21$
MaxSup $77.69\pm 0.07$
Logit Penalty $76.73\pm 0.10$
Logit Normalization [3*] 74.32
Confidence Penalty $76.58\pm 0.12$
Adaptive Label Smoothing $77.16\pm 0.15$
@@
To illustrate the differences between Max Suppression Regularization and Label Smoothing, we follow the projection technique of muller2019does. Specifically, we select three semantically related classes and construct an orthonormal basis for the plane intersecting their class templates in feature space. We then project each sample’s penultimate-layer activation vector onto this plane. To ensure the visual clarity of the resulting plots, we randomly sample 80 images from the training or validation set for each of the three classes.
@@
    Confusion. We identify a class that the Label Smoothing (LS)–trained model frequently misclassifies and select two additional classes involved in those misclassifications (Figure˜3(c), Figure˜4(c)). Conversely, we also examine a scenario where a class under Max Suppression is confused with others, highlighting key differences (Figure˜3(d), Figure˜4(d)).
@@
As shown in Figures˜4 and 3, models trained with Max Suppression exhibit:
@@
Appendix K Ablation on the Weight Schedule
We conducted an ablation study on the $\alpha$ schedule, as shown in Table˜15. The consistently high accuracy across settings demonstrates the robustness of MaxSup. The adaptive $\alpha$ schedule, adopted from [lee2022adaptive], further highlights the method’s integrity and compatibility with principled design choices.
@@
Appendix L Analysis of Computation Efficiency
@@
Since $K$ (the number of classes) is usually small (e.g., 1000 for ImageNet-1K), this overhead is minimal compared to the overall forward/backward pass of a deep network. In Table˜16, we report the average training time per epoch using a ResNet-50 model on the ImageNet-1K dataset.
@@
Appendix M Logits Visualization
As mentioned in Section˜4.1.2, all Label Smoothing variants apply a different penalty on the logits. To illustrate the impact of different methods on the logits, we plot the histogram of logits of ResNet-50 networks trained with each method over the ImageNet validation set, as shown in Figure˜5.